# A quasi-one-dimensional hydrogen-bonded monolayer ice phase

### v1

@@ Acknowledgements.
We thank Jinggang Lan, Philipp Schienbein, Michele Ceriotti, and all of A.M.’s research group members for their comments on the manuscript. P.R. would also like to thank David R. Reichman for his academic support during this project. V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship and the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge. A.M. and X.R.A. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant no. 101071937). P.R. acknowledges that: “This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0024386.” P.R. would also like to thank The Winston Churchill Foundation of the United States for their financial support. We are grateful for computational support from the Swiss National Supercomputing Centre under project s1209, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and the Cambridge Service for Data Driven Discovery (CSD3).
Appendix A Computational details
A.1 Simulation model of nanoconfined water
@@
The system considered in this work is a set of water molecules trapped between two parallel confining sheets, as shown in Fig. 5. The two parallel confining sheets are kept 5.0 Å apart, measured as the distance between the centres of the atomic nuclei in each confining sheet. This results in the water molecules forming a single monolayer parallel to each confining sheet.
@@
Figure 5: Schematic overview of the confined system and the directions of the lateral pressure and confining potential.
@@
The total potential energy of this system is written as a sum of the potential energy of the water molecules and a confining potential [8, 47, 12, 14, 10, 11, 15, 16]. The confining potential is a Morse potential fitted to Quantum Monte Carlo water-carbon total energies used in Refs. 14, 63, 16. The confining potential is uniform in the plane of confinement and depends explicitly on the perpendicular distance of a water molecule from the confining sheets. The total potential energy of the water molecules is represented using a high-dimensional Behler-Parrinello neural network potential [18] fitted to total energies and forces calculated at revPBE0-D3 density functional theory level [37, 38].
A.2 Machine Learning Potential
@@
To train the machine learning potential, we employ an active learning protocol called query by committee (QbC) implemented by  Schran et al. [19]. In the QbC approach, the MLP committee is initially trained on a small set of structures. Subsequent structures are added to the training set from a larger dictionary only if they have large committee disagreements. By only adding structures with large committee disagreements to our training set, we can iteratively and systematically build up a compact training set of informative structures. We use the bulk and interfacial water potential from Ref. 19 as a starting point and supplement additional structures using monolayer and bulk structures at a variety of temperatures and lateral pressures [16]. We added monolayer, bilayer, trilayer, and tetralayer structures from simulations of confining widths of 6.5, 9.0, 9.5, and 11.5 Å. The training error of the resulting MLP used in this work is 83.82 meV/Å for forces and 3.38 meV/water for energies. Moreover, the MLP accurately reproduces the monolayer water phase diagram [64] by  Kapil et al. [16].
A.3 vdW pressure
@@
An important component of this system is the vdW pressure (or lateral pressure) experienced between the sheets of the confining material. For instance, if the confining material is graphene, the attractive vdW forces between the two graphene sheets will pull them closer together. If a pocket of water is encapsulated between these two sheets, the graphene sheets will begin to close inwards, pushing the water pocket into an increasingly small region of space. This will continue until the internal pressure in the water balances this vdW attractive force. This results in an effective lateral pressure from the termination of the pocket that acts in the directions perpendicular to the confining potential [42]. We model this lateral pressure by running simulations in the $NP_{xy}T$ ensemble.
A.4 Molecular dynamics simulations
@@
A.5 Path integral molecular dynamics simulations
@@
The path integral molecular dynamics (PIMD) simulations used the same software and parameters as the molecular dynamics simulations. We employed simulation cells containing $144$ molecules with 32 imaginary time slices. We used an optimally-damped generalized Langevin equation thermostat to control the barostat temperature and a local path-integral Langevin equation thermostat [73] to thermalize the physical degrees of freedom. Using a BAOAB splitting of the path-integral isothermal isobaric Liouville operator, we used a timestep of 0.5 fs to run simulations for 1 ns.
Appendix B Hydrogen bond and molecular orientation definitions
B.1 Hydrogen bond definition
@@
We employ the hydrogen bond definition of  Luzar and Chandler [35]. Two water molecules are considered to be (putatively) hydrogen-bonded if they satisfy the following two conditions: (1) The distance between the two oxygen atoms is less than 3.5 Å, and (2) the angle formed by the O-O vector and one of the O-H bond vectors is less than $30^{\circ}$. The sensitivity of the hydrogen bonding trends for the flat-rhombic phase with the definition parameters are shown in Fig. 6.
@@
Figure 6: The number of (putative) hydrogen bonds as a function of temperature in the flat-rhombic phase in our classical simulations. The main curve is the same as in the classical (red) curve of Fig. 3(a). The upper/lower error bounds are obtained by increasing/decreasing the two hydrogen bonding criteria in the text by 5% each.
B.2 Molecular orientation characterization
To characterize the rotational motion of a molecule, we define $\phi$ and $\theta$ angles for molecular dipoles and a $\sigma$ order parameter for the system. The $\phi=\frac{\pi}{2}-\text{arccos}\left(\hat{\mu}\cdot\hat{z}\right)$ angle captures the alignment between the molecular dipole vector $\mathbf{\mu}$ and the $z$-axis, with hats above vectors indicating normalized unit vectors. A water molecule with $\phi=\pm\pi/2$ will have its dipole vector aligned with the $\pm$ $z$-axis. We also wrap around the $\phi$ angles computed such that $\phi$ always lies in the range $[-\frac{\pi}{2},+\frac{\pi}{2}]$. The $\theta=\text{arctan}\left({\overrightarrow{c}\cdot\hat{z}}/{\overrightarrow{c}\cdot\hat{x}}\right)$ angle captures a molecule’s rotation about its centre of mass around the $y$ axis with $\overrightarrow{c}=\overrightarrow{r_{OH_{1}}}\times\overrightarrow{r_{OH_{2}}}$ and $\overrightarrow{r_{OH_{1}}}$ and $\overrightarrow{r_{OH_{2}}}$ being the OH bond vectors of an individual water molecule. $\theta$ measures the rotation of $\overrightarrow{c}$ around the $y$-axis.
@@ B.2 Molecular orientation characterization
The $\sigma$ parameter describes a collective configuration of molecules in the flat-rhombic phase to distinguish the two states shown in Fig. 4(a). The side views of Fig. 4(a) show that the individual dipole directions of each row of water molecules change when the flat-rhombic phase exchanges between these two states. We use $\phi_{ij}$ to denote the $\phi$ angle of the water molecule in row $i$ and column $j$ of the flat-rhombic lattice. We define the switching parameter $b_{i}$ that alternates signs between rows and the $\sigma$ order parameter as
@@
where $n$ is the number of water molecules in our system. The above equation sums over the product of the row-specific $b_{i}$ parameter and the molecule-specific $\phi_{ij}$ angle between the geometric dipole of the water molecule and the confinement plane. With these definitions, the $b_{i}\phi_{ij}$ terms in the summation will have the same sign if the rows of water molecules have alternating dipole directions. Therefore, the absolute value of this order parameter $|\sigma|$ should take on large values for the crystal structure of the flat-rhombic phase, and the motion shown in Fig. 4(a) should correspond to a sign change in this $\sigma$ order parameter. Given the arbitrary choice of positive values for odd $i$ and negative values for even $i$ in the definition of $b_{i}$, only relative changes in the sign of $\sigma$ should be considered, as the absolute sign of $\sigma$ has no meaning.
Appendix C Hydrogen bonding characterization of the hexagonal and pentagonal phases

### v2

@@ Acknowledgements.
We thank Jinggang Lan, Philipp Schienbein, Michele Ceriotti, and all of A.M.’s research group members for their comments on the manuscript. P.R. would also like to thank David R. Reichman for his academic support during this project. V.K. acknowledges support from the Ernest Oppenheimer Early Career Fellowship and the Sydney Harvey Junior Research Fellowship, Churchill College, University of Cambridge. A.M. and X.R.A. acknowledge support from the European Union under the “n-AQUA” European Research Council project (Grant no. 101071937). C.S. acknowledges partial financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) project number 500244608. P.R. would like to thank The Winston Churchill Foundation of the United States for their financial support. P.R. would also like to acknowledge that: “This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0024386. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.” We are grateful for computational support from the Swiss National Supercomputing Centre under project s1209, the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and the Cambridge Service for Data Driven Discovery (CSD3).
Author Contributions
P.R., A.M., and V.K. conceived the study. P.R., V.K., and C.S. trained the machine learning potential used in this work. P.R. and X.R.A. performed molecular dynamics simulations using this MLP. All of the authors were involved in discussing results and assembling the manuscript.
Competing Interests
The authors declare no competing interests.
Figure Legends
Fig. 1 |||| Nanoconfined ice phases and the role of vdW stacking of hydrogen bonded chains. a) The four nanoconfined ice phases considered in this work from Ref. 8 and Ref. 14, along with the lateral pressures used for simulating each phase. Solid lines between water molecules indicate hydrogen bonds. We highlight chains of hydrogen-bonded water molecules in each phase. For estimating $E_{\mathrm{chain}}$, we use the unique highlighted chains in the case of the flat-rhombic and ZZ-qBI phases. For the hexagonal and pentagonal phases, we report the average value and the standard deviation as the error bar, where the standard deviation is computed across different choices of chains, as described in the text. b) For each phase, we compute the stabilization energy of a single molecule in the 0 K crystal structure, $E_{\mathrm{lattice}}$; the (average) stabilization energy of a water molecule within the hydrogen bonded chain(s), $E_{\mathrm{chain}}$; and the stabilization energy between chains in the crystal structure, $E_{\mathrm{stack}}$. The hatched bars show the contribution of vdW interactions to the corresponding stabilization energy. Source data are provided as a Source Data file.
Fig. 2 |||| Depiction of the orientations of water molecules. a) The $\phi$ angle captures in-plane and out-of-plane fluctuations of water molecules’ dipole moments. b) The $\theta$ angle captures in-plane rotations of water molecules. We refer the readers to section II.B for the mathematical definitions of the $\phi$ and $\theta$ angles. c) The left and right images depict the molecular orientations at the probability maxima in Fig. 3(a), while the middle image depicts the molecular orientation at the maximum for the number of putative hydrogen bonds in Fig. 3(b).
Fig. 3 |||| Molecular orientation and the number of putative hydrogen bonds. a) The two-dimensional log probabilities along the $\phi$ and $\theta$ parameters for the flat-rhombic phase at 2 GPa and three different temperatures. b) The average number of putative hydrogen bonds associated with each $(\theta,\phi)$ molecular orientation for the flat-rhombic phase under the same conditions. Relevant molecular orientations are depicted in Fig. 2(c). We employ the geometric definition of the hydrogen bond from Ref. 29. Source data are provided as a Source Data file.
Fig. 4 |||| Temperature dependence of hydrogen bonding. a) The average number of putative/geometric hydrogen bonds across the range of temperatures in which the flat-rhombic phase is stable using classical and quantum simulations. The error bars show the standard error of the mean, as computed by block averaging over 10 blocks. b) The proportion of water molecules of each hydrogen bonding motif (defined in the text). The 1D2A proportions are perfectly aligned with the 2D1A proportions. c) The 2D1A motif’s lifetime remains extremely short across the range of temperatures in which it is observed. The error bars show the standard error of the mean, computed across all instances of hydrogen bonding. Dashed lines serve as a visual guide for the eye. Source data are provided as a Source Data file.
Fig. 5 |||| Temperature dependence of the dielectric response. The classical dielectric constant of the flat-rhombic phase across the range of temperatures considered in this work. The singularity at 380 K corresponds to the transition to the hexatic phase of nanoconfined water [15]. Source data are provided as a Source Data file.
Fig. 6 |||| Hydrogen bonded switching behavior. a) The two symmetry-related structures of the flat-rhombic phase, with alternating rows of aligned dipoles, discerned by the different signs of $\sigma$. A zero value of $\sigma$ indicates that the molecular dipoles are not aligned row-wise. The side views in each panel show the system as viewed from the right side, where the switch in the row-dependent dipole direction between the two states is clear. b) For smaller unit cells with 144 water molecules, we observe coherent switching between these two states at intermediate temperatures, as shown in the $\sigma$ trajectories for three different temperatures. c) The free energy profiles for this $\sigma$ parameter show that raising the temperature lowers the free energy barrier between these states. The smooth, low-barrier free energy profile at high temperatures indicates that the molecular dipoles lose spatial coherence. Source data are provided as a Source Data file.
Supplementary Information:
Quasi-one-dimensional hydrogen bonding in nanoconfined ice
I Computational Details
I.1 Simulation model of nanoconfined water
@@
The system considered in this work is a set of water molecules trapped between two parallel confining sheets, as shown in Supplementary Figure 7. The two parallel confining sheets are kept 5.0 Å apart, measured as the distance between the centers of the atomic nuclei in each confining sheet. This results in the water molecules forming a single monolayer parallel to each confining sheet.
@@
Fig. 7: |||| Simulation setup overview. Schematic overview of the confined system and the directions of the lateral pressure and confining potential.
@@
The total potential energy of this system is written as a sum of the potential energy of the water molecules and a confining potential [1, 2, 3, 4, 5, 6, 7, 8]. The confining potential is a Morse potential fitted to Quantum Monte Carlo water-carbon total energies used in Refs. 4, 9, 8. The confining potential is uniform in the plane of confinement and depends explicitly on the perpendicular distance of a water molecule from the confining sheets. The total potential energy of the water molecules is represented using a high-dimensional Behler-Parrinello neural network potential [10] fitted to total energies and forces calculated at revPBE0-D3 density functional theory level [11, 12].
I.2 Machine Learning Potential
@@
To train the machine learning potential, we employ an active learning protocol called query by committee (QbC) implemented by Schran et al. [13]. In the QbC approach, the MLP committee is initially trained on a small set of structures. Subsequent structures are added to the training set from a larger dictionary only if they have large committee disagreements. By only adding structures with large committee disagreements to our training set, we can iteratively and systematically build up a compact training set of informative structures. Previous work used the bulk and interfacial water potential from Ref. 13 as a starting point and supplemented additional monolayer and bulk structures at a variety of temperatures and lateral pressures [8]. In this work, we further added monolayer, bilayer, trilayer, and tetralayer structures from simulations of confining widths of 6.5, 9.0, 9.5, and 11.5 Å. We also included structures from PIMD simulation trajectories so that the training dataset includes the new configurations that are accessible with the inclusion of nuclear quantum effects. The average training error of the resulting MLP used in this work is 83.82 meV/Å for forces and 3.38 meV/water for energies.
To assess the neural network’s ability to generalize to unseen configurations, we compared the energies and forces produced by the MLP to those produced by the reference revPBE0-D3 functional on new structures that were not seen during training time. These structures were generated by running 2 nanosecond $NP_{xy}T$ simulations at 2.0 GPa starting from each of the solid phases identified in Ref. 8. Each solid phase was simulated at 100 K, 300 K, and 600 K. For some of these simulations, the solid phase was metastable, and so the trajectory remained in the solid phase. For other simulations, the solid phase was unstable, and the water molecules in the simulation entered a disordered phase or glassy phase (depending on the conditions). Hence, these trajectories contained a diverse set of ordered and disordered configurations that span a broad range of local molecular configurations. We computed the errors between the MLP’s predictions and the reference revPBE0-D3 calculations over all of the configurations generated in each of these trajectories. The resulting mean absolute errors (MAE) for the energies and root-mean-squared errors (RMSE) for average atomic forces are shown in Supplementary Table 2. These errors are similar in magnitude to the errors of MLPs employed in other studies [8, 14], which have been validated against reference DFT simulations. We also note that this model [15] (and its predecessors based on the same machine learning architecture [16, 17]) have been validated against experimental IR, Raman and sum frequency generation spectra, demonstrating an excellent description of real-time dynamics of aqueous systems.
Table 2: |||| Force field validation. Average MLP errors on configurations that were not seen during training time. These configurations comes from $NP_{xy}T$ simulations at 2.0 GPa covering a wide range of newly-generated ordered and disordered structures.
Energy MAE Avg Atomic Force
Temperature [meV/water] RMSE [meV/Å]
100 K 5.8 67.6
300 K 4.8 81.8
600 K 5.6 103.4
I.3 vdW pressure
@@
An important component of this system is the vdW pressure (or lateral pressure) experienced between the sheets of the confining material. For instance, if the confining material is graphene, the attractive vdW forces between the two graphene sheets will pull them closer together. If a pocket of water is encapsulated between these two sheets, the graphene sheets will begin to close inwards, pushing the water pocket into an increasingly small region of space. This will continue until the internal pressure of the water balances this attractive vdW force. This results in an effective lateral pressure from the termination of the pocket that acts in the directions perpendicular to the confining potential [18]. We model this lateral pressure by running simulations in the $NP_{xy}T$ ensemble.
I.4 Molecular dynamics simulations
@@
I.5 Confining Potential
As discussed earlier, we used an implicit confining potential to emulate the interactions between the water molecules and the confining material. The two parallel walls are kept at a fixed distance of 5 Å away from each other, and the nanoconfined water molecules all lie between these two implicit walls. If a water molecule is a distance $z$ away from one of the walls, the potential that it experiences is a Morse potential of the form:
$V_{morse}=D_{0}\quantity[\quantity(1-e^{-a\quantity(z-z_{0})})^{2}-1]$ (1)
with the parameter values $D_{0}=5.78\crossproduct 10^{-2}$ eV, $z_{0}=3.85$, and $a=0.92\text{ \AA{}}^{-1}$. These values were obtained from a fit to water-carbon Quantum Monte Carlo (QMC) interaction energies [27].
I.6 Path integral molecular dynamics simulations
@@
The path integral molecular dynamics (PIMD) simulations used the same software and parameters as the molecular dynamics simulations. We employed simulation cells containing $144$ molecules with 32 imaginary time slices. We used an optimally-damped generalized Langevin equation thermostat to control the barostat temperature and a local path-integral Langevin equation thermostat [28] to thermalize the physical degrees of freedom. Using a BAOAB splitting of the path-integral isothermal isobaric Liouville operator, we used a timestep of 0.5 fs to run simulations for 1 ns. For structural and thermodynamic quantities (e.g., free energy profiles), we average results over all the imaginary time slices from the PIMD simulations. For dynamical quantities, we use the centroids of the imaginary time slices.
I.7 Computing free energy profiles
We used a simple binning procedure to compute free energy profiles along a particular variable, $x$. Given a molecular dynamics trajectory, we assemble a list of all values that the variable $x$ takes on in each frame of the trajectory. We then compute the probability histogram $P(x)$ with an appropriately chosen number of bins. The free energy profile in units of $k_{B}T$ is then reported as $F(x)=-\log(P(x))$, where $k_{B}$ is Boltzmann’s constant. To account for quantum nuclear effects, we perform the binning procedure on the trajectory of the individual imaginary time slices of our PIMD simulations and average the histograms.
II Hydrogen Bond Definition
@@
We employ the geometric hydrogen bond definition of Luzar and Chandler [29]. Two water molecules are considered to be (putatively) hydrogen-bonded if they satisfy the following two conditions: (1) The distance between the two oxygen atoms is less than 3.5 Å, and (2) the angle formed by the O-O vector and one of the O-H bond vectors is less than $30^{\circ}$. The sensitivity of the hydrogen bonding trends for the flat-rhombic phase with respect to the parameters of this geometric definition are shown in Supplementary Figure 8.
In Fig. 4(a) of the main text, we showed that employing such a geometric hydrogen bond definition resulted in the number of hydrogen bonds seeming to increase as temperature is raised. This contradicts the expected behavior from a standard energy-entropy trade-off perspective. We suggested that this was because this geometric definition still included fleeting hydrogen bonds that only lasted for times shorter than intermolecular timescales. Employing the definition by Schienbein and Marx [30], we also computed the number of dynamical hydrogen bonds as a function of temperature. Here, a dynamical hydrogen bond is one that satisfies the above geometric criteria for at least 0.2 ps, the typical timescale for intermolecular oscillations for monolayer ice as well as supercritical and room temperature water [30]. The resulting dynamical hydrogen bond count exhibits the expected decreasing trend, as shown in Supplementary Figure 9.
@@
Fig. 8: |||| Geometric hydrogen bond definition. The number of (putative) hydrogen bonds in our classical simulations of the flat-rhombic phase as a function of temperature. The main curve is the same as in the classical (red) curve of Fig. 4(a) in the main text. The upper/lower error bounds are obtained by increasing/decreasing the two hydrogen bonding criteria in the text by 5% each.
Fig. 9: |||| Dynamical hydrogen bond count. The grey data points show the number of hydrogen bonds that survive at least 0.2 picoseconds, with the lower and upper error bounds given by a 0.3 and 0.1 picosecond cutoff, respectively. The red and black lines are the same as in Fig. 4(a) of the main text, showing the number of geometric/putative hydrogen bonds as a function of temperature. The error bars for the putative lines show the standard error of the mean, as computed by block averaging over 10 blocks.
III $\sigma$ Order Parameter Definition
@@
To describe the concerted motion shown in Fig. 6(a) of the main text, we define a $\sigma$ order parameter to distinguish the two states. The side views of Fig. 6(a) show that the individual dipole directions of each row of water molecules change when the flat-rhombic phase exchanges between these states. We use $\phi_{ij}$ to denote the $\phi$ angle of the water molecule in row $i$ and column $j$ of the flat-rhombic lattice. We define the switching parameter $b_{i}$ that alternates signs between rows and the $\sigma$ order parameter as
@@
where $n$ is the number of water molecules in our system. The above equation sums over the product of the row-specific $b_{i}$ parameter and the molecule-specific $\phi_{ij}$ angle between the geometric dipole of the water molecule and the confinement plane. With these definitions, the $b_{i}\phi_{ij}$ terms in the summation will have the same sign if the rows of water molecules have alternating dipole directions. Therefore, the absolute value of this order parameter $|\sigma|$ should take on large values for the crystal structure of the flat-rhombic phase, and the motion shown in Fig. 6(a) should correspond to a sign change in this $\sigma$ order parameter. Given the arbitrary choice of positive values for odd $i$ and negative values for even $i$ in the definition of $b_{i}$, only relative changes in the sign of $\sigma$ should be considered, as the absolute sign of $\sigma$ has no meaning. Furthermore, note that during the concerted motion described above, the value of the $\sigma$ order parameter will switch its sign very quickly. If the motion of the protons is not correlated, then the values of the $\phi_{ij}$ values will be uncorrelated with the row index $i$. This will cause the $\sigma$ value to take on intermediate values near 0, which is indicative of a loss of proton order.
IV Rotational Autocorrelation Functions
Fig. 10: |||| Rotational Legendre Autocorrelation Functions. a) The rotational Legendre autocorrelation functions of water molecules in the flat-rhombic phase at various temperatures. These are computed as described in Ref. [31] with $n=1$. These are averaged over 5 different trajectories. The width of each line is the error, as computed by the standard error of the mean across the trajectories. b) The temperature dependence of the rotational relaxation time $\tau$, fit to an Arrhenius form (dashed black line). We estimate $\tau$ as the integral of the above autocorrelation functions, and we only consider temperatures where the concerted motion is observed. The error bars for $\ln(\tau)$ are computed by propagating the errors from the top panel.
For dynamical insights into the rotational motion of water molecules, we estimate the rotational Legendre autocorrelation functions of the O–H bond with $n=1$, following the procedure described in Ref. [31]. This autocorrelation function will decay to 0 as the initial rotational configuration of the water molecule relaxes to a random orientation. For each displayed temperature in Supplementary Figure 10(a), we estimate the autocorrelation function over a 25 picosecond window, averaged over 5 trajectories. The definite integral of this function, from 00 to $\infty$, gives the rotational relaxation time.
Supplementary Figure 10(a) clearly shows two different types of rotational motion in the flat-rhombic phase. As we mentioned in the discussion surrounding Fig. 3 of the main text, water molecules in low-temperature conditions remain localized in their molecular orientations up to the simulation time considered in this work. At low temperatures such as 160 and 220 K, this behaviour leads to a sharp initial decay in the rotational Legendre autocorrelation plots in Supplementary Figure 10(a). This sharp decay is due to thermal fluctuations within the $(\theta,\phi)$ free energy minimum. As a result, these autocorrelation functions exhibit a nearly infinite rotational relaxation time, since the water molecules are stuck in the free energy minimum that they begin in. The diverging relaxation time is an artefact of the absence of molecular rotation sampling within the timescales of our simulations. At intermediate temperatures such as 280 K and 360 K, we observe finite rotational relaxation times, consistent with the $(\theta,\phi)$ exploration of water molecules in Fig. 3 of the main text. We also note a good fit of the rotational relaxation times between 280 K and 360 K to an Arrhenius equation, suggesting a common activation barrier dictating the rotational motion. We do not include the temperatures 160 and 220 K in the fit, as at these temperatures we do not sample the relevant relaxation event, implying an $\infty$ relaxation time. We do not account for quantum nuclear motion as they are expected to only make a quantitative difference to this trend [32, 33]. Specifically in the context of the flat-rhombic phase, the correspondence between Fig. 3 of the main text and Supplementary Figure 10, and the quantitative increase in the exploration of $(\theta,\phi)$ orientations due to quantum nuclear effects (see Supplementary Figure 13) implies a systematic increase the relaxation time.
V Hydrogen Bonding Characterization of the Hexagonal and Pentagonal Phases